Suppose ACT Mathematics scores are normally distributed with a mean of 20.8 and a standard deviation of 5.4. A university plans to admit students whose scores are in the top 30%. What is the minimum score required for admission?
Answer by Boreal(15235)

z from (-oo,+0.525)=0.700.
Therefore, any score greater than 0.525 sd from the mean will be acceptable.
20.8+5.4(0.525)=20.8+2.835=23.64 would be the minimum.